NO HANDWRITTEN ANSWERS PLEASE

Inspired by his trip to Greece, the gentleman farmer decided he needed a Mykonos-style windmill pumping a pond for his sheep and orchard. His earned value table for the project appears below. At the end of June, what were the estimated time to completion and the estimated cumulative cost to completion?

Activity

Jan

Feb

Mar

Apr

May

Jun

Jul

Plan

% Complete

Earned Value

Grading

500

150

650

100%

650

Footings

200

150

350

100%

350

Foundation

800

300

1,100

80%

880

Masonry

300

300

200

800

70%

560

Interior

700

2,000

2,700

50%

1,350

Pump Install

3,000

3,000

40%

1,200

Painting

350

400

750

25%

188

Monthly Plan

500

350

950

600

1,000

2,550

3,400

9,350

5,178

Cumulative

500

850

1,800

2,400

3,400

5,950

9,350

Monthly Actual

600

400

900

800

800

2,800

0

Cumulative Actual

600

1,000

1,900

2,700

3,500

6,300
Answer #1

Cost Performance Index (CPI) = EV / AC = 5,178 / 6,300 = 0.8219 (overbidget)
Schedule Performance Index (SPI) = EV / PV = 5,178 / 5,950 = 0.8703 (delay)

Assuming that the same sort of delay will happen in rest of the project, the estimated time to complete the remaining project is:

= 1 month / SPI

= 1 / 0.873

= 1.15 months

So, total time for the project = 7.15 months.

The estimated cumulative cost to completion = Estimate At Completion (EAC)

= AC + (BAC - EV) / (SPI * CPI)

= 6300 + (9350 - 5178) / (0.8219 * 0.8703)

= $12,133